### v1

@@ 5.2 Interpretation from ML paradigms
As we can see from the stated goal in the application domain, the criteria for success of the resulting model is to measure its predictive performance against empirical results. Mapping between input and output pairs is non-trivial in a lot of cases. The information obtained through experimental measurements does not always align directly with what LGEM+ can predict, particularly with relation to current metabolomics methods, However, despite the imperfect mapping of predicted to observed outputs, this follows the supervised learning paradigm.
With regards to active learning, we recall that exploration of the experimental space was due to uncertainty of the model in particular neighbourhoods. Due to the complexity of the system and the stochastic factors, certain phenomena might require hundreds of experiments which are very close together in the experimental space. This requires precision, and Genesis has been designed to provide this precision.
@@ 6 Conclusions and Future Directions
One motivation for choosing FOL to model yeast metabolism was grounded in epistemic parsimony; FOL frameworks are easily extensible. There are many phenomena that could be relevant for experimentation, that LGEM+ does not currently model. But we can update the logical framework without affecting the underlying infrastructure. We saw from our results on the single-gene essentiality prediction that there is a gap in explanation. We need to accede a more complex model, and the background knowledge suggests that we should incorporate new mechanisms such as gene regulation to capture some of the higher order behaviours of the system.
Our next steps will be working toward full integration of LGEM+ with the Genesis hardware, relying on controlled vocabularies [26] to specify experiments and results so that we can map inputs to outputs for the machine learning algorithms. These are necessary steps so that we can employ LGEM+ and Genesis for closed-loop experimentation.
Acknowledgements

### v2

@@ 5.2 Interpretation from ML paradigms
As we can see from the stated goal in the application domain, the criteria for success of the resulting model is to measure its predictive performance against empirical results. Mapping between input and output pairs is non-trivial in a lot of cases. The information obtained through experimental measurements does not always align directly with what LGEM+ can predict, particularly with relation to current metabolomics methods. However, despite the imperfect mapping of predicted to observed outputs, this follows the supervised learning paradigm.
Models like LGEM+ provide plausible hypothesis candidates for testing and feeding back to the discovery loop, which is aligned with the active learning paradigm and targeted exploration. When exploring the experimental space, due to the complexity of the system and the stochastic factors, certain phenomena might require many experiments that are very close together in the experimental space. This requires precision, and Genesis has been designed to provide this precision. The other selection criterion was to increase diversity in the dataset. A general purpose microformulator and modular design were deliberately chosen to increase options for experimentation.
@@ 6 Conclusions and Future Directions
One motivation for choosing FOL to model yeast metabolism was grounded in epistemic parsimony; FOL frameworks are easily extensible. There are many phenomena that could be relevant for experimentation, that LGEM+ does not currently model. But the logical framework can be updated without affecting the underlying infrastructure. We saw from the results on the single-gene essentiality prediction that there is a gap in explanation. It is necessary to to accede a more complex model, and the background knowledge suggests that we should incorporate new mechanisms such as gene regulation to capture some of the higher order behaviours of the system.
Our next steps will be working toward full integration of the computational models with the Genesis hardware, relying on controlled vocabularies [31] to specify experiments and results so that we can map inputs to outputs for the machine learning algorithms. These are necessary steps so that we can employ computational models like LGEM+ with Genesis for closed-loop experimentation.
{credits}
6.0.1 Acknowledgements